OpenStudy (anonymous):

If f′ is continuous, f(7)=0, and f′(7)=11, evaluate

OpenStudy (dan815):

?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} \frac{ f(7+3x)+f(7+5x)}{ x}\]

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

ok

OpenStudy (dan815):

lhos[itsl since inderminate form

OpenStudy (dan815):

l'hopital*

OpenStudy (aum):

\[ \lim_{x \rightarrow 0} \frac{ f(7+3x)+f(7+5x)}{ x} = \lim_{x \rightarrow 0} \frac{ f(7+3x)}{x} +\lim_{x \rightarrow 0} \frac{(7+5x)}{ x} = \\ \lim_{x \rightarrow 0} 3\frac{ f(7+3x)}{3x} +\lim_{x \rightarrow 0} 7\frac{(7+5x)}{ 7x} \]

OpenStudy (aum):

= 3f'(7) + 7f'(7)

OpenStudy (dan815):